### v1

# Investigating MAXI J1752--457 with NuSTAR in the aftermath of a superburst

Abstract
We present two NuSTAR observations of the X-ray transient, MAXI J1752--457, following a superburst which was observed by MAXI/GSC in November, 2024. NuSTAR follow-up confirmed that MAXI J1752--457 is coincident with the previously observed Einstein Probe source, EP240809a. We performed a spectral analysis of the source during both NuSTAR observations, and we confirm that the hard X-ray spectra are consistent with the inclusion of a spherical blackbody component. At about 79 hours after the onset of the superburst, we find a blackbody temperature of $kT_{\mathrm{bb}}=0.60\pm 0.1$ keV and $R_{\mathrm{bb}}/D_{8}=6.0^{+0.4}_{-0.3}$ km (not including corrections for scattering in the neutron star atmosphere), where $D_{8}$ is the source distance, which is not yet known, in units of 8 kpc. Furthermore, we report a hard X-ray excess which is fit well by a power law with photon index $\Gamma\approx 4$, much steeper than those typically seen during accretion onto neutron stars at similar luminosities. We infer that the electron energy distribution in the Comptonizing medium which gives rise to the power law component differs significantly compared to purely accretion-powered neutron star transients even several days after the onset and rapid decay of superbursts. We also performed an energy-resolved timing analysis which showed that the source variability was dominated by red noise in the Comptonization component, suggesting coupling with an accretion disk, rather than being seeded by thermal emission from the neutron star surface.
Neutron stars — thermonuclear bursts — X-ray transients
facilities: MAXI/GSC, NuSTARsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024a, b), Xspec (Arnaud, 1996), SAOImageDS9 (Joye & Mandel, 2003)
1 Introduction
X-ray binaries (XRB), formed via accretion onto a compact object from a companion, exhibit a wide array of variability on timescales ranging from milliseconds to days to months. In particular, most XRBs are transient, meaning that they alternate between intervals of quiescence and outburst due to changes in the rate of mass accretion. For example, in Be X-ray binaries consisting of a Be star and an accreting neutron star, these periods of outburst are caused by transient accretion when the compact object passes near its companion in its orbit. XRBs hosting black holes undergo weeks- or months-long outbursts due to slow accumulation of material into their accretion disks, leading to a critical point at which material accretes much more efficiently onto the compact object (Mineshige & Wheeler, 1989).
These accretion-powered outbursts are distinct from bursting behavior caused by thermonuclear burning. Neutron star XRBs which persistently accrete from low mass companions at a relatively low level may accumulate material on their surfaces without burning until a critical depth is reached, producing runaway burning of accreted hydrogen and helium. We observe this runaway conflagration in the form of a rapid flash known as a Type I X-ray burst, lasting only seconds to minutes, during which the source luminosity may increase by multiple orders of magnitude up to the Eddington luminosity. The depth at which accumulated material ignites plays a decisive role in the duration and structure of the resulting burst. Some sources exhibit “superbursts” which resemble Type I X-ray bursts but which last several hours and decay on the timescale of days due to ignition at much larger column depths producing extended carbon burning (in’t Zand, 2017; Cumming, 2003; Keek et al., 2008; Serino et al., 2016). Due to their extended timescale, superbursts can provide an opportunity to probe how accretion disks and coronae respond to high levels of X-ray irradiation. The rarity of these events, however, has thus far limited investigations into superbursts. In this paper, we present an analysis of the newly-discovered source, MAXI J1752--457, which was observed by the Nuclear Spectroscopic Telescope Array (NuSTAR) during the decay phase of a superburst.
The first detection of MAXI J1752--457 with the Monitor of All-sky X-ray Image (MAXI) Gas Slit Camera (GSC) occurred during the scan transit at 18:23 on November 9, 2024 UTC (MJD 60623.766) with a 4–10 keV flux of $982\pm 50$ mCrab, or $(1.12\pm 0.06)\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$(Serino et al., 2024). The flux dropped rapidly over subsequent scan transits. Several months prior to the detection of MAXI J1752--457 with MAXI, an X-ray transient, deemed EP240809a, was detected by Einstein Probe (EP) with a position consistent with MAXI J1752--457 (Liu et al., 2024). This earlier period of brightening, which was observed on August 9, 2024 (UTC), reached a much lower flux than the November burst: Liu et al. (2024) reported an unabsorbed 0.4–5 keV flux of $5.3\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$, corresponding to about 1.5 mCrab. An optical flare from the source was also observed by the Asteroid Terrestrial-impact Last Alert System (ATLAS) beginning about three weeks before the EP detection (Zhu et al., 2024). The authors pointed out that the soft X-ray and the blue optical spectra were consistent with a tidal disruption event (TDE), but no host galaxy was found. Therefore, the nature of the source remained unknown.
MAXI/GSC observations showed a thermal spectrum which softened during the rapid decay, leading Negoro et al. (2024) to propose that the origin of the the source activity was a long X-ray burst from an uncatalogued X-ray binary, rather than a TDE. By combining MAXI/GSC and NinjaSat observations of the November burst, Aoyama et al. (2025) have shown that the source is indeed consistent with an accreting neutron star. The flux evolution of the burst, which exhibits a fast rise and exponential decay on the timescale of days, along with an X-ray spectrum which can be described with a spherical blackbody, indicate that the transient event corresponded to accretion-induced thermonuclear burning. In particular, the hours-long timescale of the burst and the years-long recurrence time (inferred from the lack of previous detections of the source by MAXI), indicate a superburst.
At the time that it was detected by MAXI/GSC, MAXI J1752--457 had a small Solar Aspect Angle (SAA) of $<45^{\circ}$, making follow-up observations impossible for most pointing X-ray observatories. Thanks to the ability of NuSTAR to point near or at the Sun, we were able to trigger hard X-ray observations of the MAXI confidence region. These observations confirmed that MAXI J1752--457 is coincident with EP240809a, and preliminary spectral analysis provided the first evidence for a hard X-ray excess in addition to the blackbody spectral component (Pike et al., 2024).
Here, we present a detailed analysis of two NuSTAR observations of MAXI J1752--457 following its November, 2024, superburst, and we place the flux of the source as measured by NuSTAR in the context of MAXI/GSC monitoring of the burst. In Section 2, we briefly present the MAXI/GSC observations of the burst. In Section 3, we present both spectral and timing analyses of the source with NuSTAR. Finally, in Section 4, we discuss our results and conclusions. All errors are quoted at the $90\%$ confidence level unless otherwise stated.
2 MAXI light-curve
Table 1: Details of the NuSTAR observations.
OBSID Start time (MJD) End time (MJD) Exposure (ks)
81010302001 60626.32 60626.40 6.6
81001347001 60627.05 60627.55 26
Figure 1: Top: MAXI/GSC 2–10 keV count rate (blue) and 2–10 keV flux measured with NuSTAR (yellow). Flux was converted to MAXI/GSC count rate using WebPIMMS. The vertical red dashed line marks MJD 60623.766, which is the time of the first scan when the source was reported to be detected by MAXI (Serino et al., 2024). Bottom: Background subtracted NuSTAR FPMA+FPMB light-curves. NuSTAR light-curves were reconstructed from both Mode 1 and Mode 06 data.
MAXI, launched in 2009, is an all-sky X-ray monitor installed on the Japanese Experiment Module onboard the International Space Station (Matsuoka et al., 2009). It consists of two instruments: the Gas Slit Camera (GSC) with a net bandpass of 2–20 keV and field of view (FOV) of $1.5^{\circ}\times 160^{\circ}$ (Mihara et al., 2011), and the Solid-state Slit Camera (SSC) with a net bandpass of 0.7–7 keV and FOV of $1.5^{\circ}\times 90^{\circ}$. The SSC ceased operations on June 22, 2022, and was therefore not used during observations of MAXI J1752--457. MAXI/GSC scans the full X-ray sky once every orbit (about 92 minutes), such that individual sources are within the FOV for about $45$ seconds per orbit (Sugizaki et al., 2011). The resulting 3-$\sigma$ detection limit for one scan is $80$ mCrab in the 3–10 keV band (Negoro et al., 2016). The MAXI/GSC light-curve of MAXI J1752--457 available from the online MAXI light-curve portal [^1] http://maxi.riken.jp/star_data/J1752-457/J1752-457.html contained data with a low signal-to-noise ratio due to contamination by the bright low-mass XRB (LMXB), H 1735--444, separated from MAXI J1752--457 by $2.8^{\circ}$ in the GSC anode direction. Therefore, we instead obtained the MAXI/GSC light-curve using the point-spread-function fit method presented by Morii et al. (2016). We show the MAXI/GSC light-curve for MJD 60622 – MJD 60629 in the top panel of Figure 1. We have re-binned the data by up to 10 orbits for clarity before and after the brightest phase of the burst. We note that visual inspection of the MAXI/GSC light-curve of MAXI J1752--457 binned into 16 days over the course of more than 15 years reveals a hint of a weak enhancement ($\sim$ 1–2 mCrab) beginning in July, 2024, and lasting roughly through November, 2024.
3 NuSTAR Data and Analysis
NuSTAR is the first focusing hard X-ray telescope (Harrison et al., 2013). It consists of two focal plane modules, FPMA and FPMB, each of which is composed of a set of four Cadmium Zinc Telluride (CZT) detectors in a $2\times 2$ configuration. FPMA and FPMB are both paired with a focusing optics module at a focal length of 10 m. The resulting energy range of NuSTAR is 3–79 keV, and the FOV is $10^{\prime}$ at 10 keV.
MAXI J1752--457 was observed twice by NuSTAR during the tail of the 2024 superburst. The first observation was performed using a tiling strategy designed to locate the source within the MAXI/GSC 90% confidence region. Four tiling observations were carried out with exposures of about $5$ ks each. The source was detected in one of these observations with OBSID 81010302001. Using the resulting localization, a second, pointed observation (OBSID 81001347001) was carried out less than a day later. The start and end times and the total exposure times of the two NuSTAR observations are listed in Table 1. We extracted cleaned event lists for both observations using the standard nupipeline options with saamode=NONE and tentacle=no. We used HEASOFT v6.34, NuSTARDAS v2.1.4a, and CALDB v20241104.
Figure 2: Mode 06 images taken by FPMA during the first (left) and second (right) NuSTAR observations. We split Mode 06 data into different CHU combinations in order to mitigate the effects of centroid smearing. Source and background extraction regions are shown with solid and dashed lines, respectively, for CHU combinations CHU23 (cyan), CHU3 (magenta), and CHU13 (red). A region of stray light can be seen in the northwest corner of each image. We attempted to choose source regions such that stray light contamination of the science products was minimized. For FPMB, the source was further away from the region of stray light during the first observation, allowing us to use slightly larger source regions than we chose for FPMA.
During both NuSTAR observations, the source was at a low SAA, leading to complications in the image reconstruction process. At low SAA, the primary star tracker, Camera Head Unit 4 (CHU4), which is located on the X-ray optics bench, may be blinded by the Sun and therefore become unusable for image reconstruction. The usual science mode, which utilizes CHU4, is called Mode 1. Outside of Mode 1, image reconstruction must be performed using combinations of CHUs 1, 2, and 3, which are located on the spacecraft bus. This is referred to as Mode 06 data. Depending on the pointing of the spacecraft any given time, one of seven combinations of these three star trackers may be used to produce an attitude solution. Due to thermal flexing of the mountings of CHUs 1, 2, and 3, the resulting solution may include inaccuracies which manifest as a “smearing” of the source or the presence of multiple centroids in the reconstructed image. In order to most effectively extract scientific products from Mode 06 data, we used the nusplitsc command to split the cleaned event list according to active CHU combination. We performed nusplitsc two separate times, once with splitmode=NORMAL and another time with splitmode=STRICT. In some cases, strict splitting can help to remove flickering between CHU combinations, therefore mitigating multiple centroids. However, based on visual inspection using SAOImageDS9 (Joye & Mandel, 2003), we determined that using the strict splitmode did not improve the image reconstruction compared to normal mode, and in fact strict mode only resulted in a reduction in the overall number of events available for science products. We therefore proceeded using the event files produced using splitmode=NORMAL.
We used DS9 to inspect the event files from Mode 1 and the split event files from Mode 06. For each file, we chose a circular source extraction region and centered it on the source using the DS9 centroid calculation feature. During the first observation, there was significant stray light which did not overlap with the source position, but which was close enough to the source to become a consideration when choosing the source region. In order to avoid stray light contamination of the source data, we chose source regions with radius $35^{\prime\prime}$ for FPMA data and $40^{\prime\prime}$ for FPMB data from the first observation. For both FPMA and FPMB data from the second observation, we chose source regions with radius $45^{\prime\prime}$. For all event files, we chose circular background regions with radius $100^{\prime\prime}$. The stray light region, characteristic centroid smearing, and examples of the source and background regions we chose for FPMA are shown in the Mode 06 images for both observations in Figure 2. After selecting extraction regions, we used nuproducts to produce source and background spectra and light-curves for each cleaned event file. The background-subtracted light-curves (summed FPMA and FPMB) are shown in the bottom panels of Figure 1. Finally, for each observation, we used addspec to sum the Mode 1 and Mode 06 data into one pair of FPMA source and background spectra and one pair of FPMB spectra. We used ftgrouppha to rebin the spectra according to the method described by Kaastra & Bleeker (2016) with a minimum of 10 counts per bin.
3.1 Spectral Analysis
Figure 3: NuSTAR FPMA (blue) and FPMB (orange) spectra of the source during the first (left) and second (right) observations. The dashed and dotted lines show the blackbody and power law components of the spectral model, respectively.
Table 2: Spectral parameters and unabsorbed 2–10 keV flux measured during the NuSTAR observations.
Model Component Parameter Obs1 value Obs2 value
bbodyrad $kT\,\mathrm{(keV)}$ $0.61\pm 0.02$ $0.60\pm 0.01$
$R\,\mathrm{(km)}$$\dagger$\$$\dagger$\$footnotemark: $7.7^{+1.2}_{-1.4}$ $6.0^{+0.4}_{-0.3}$
powerlaw $\Gamma$ $4.0\pm 0.4$ $3.7\pm 0.2$
$\chi^{2}/\mathrm{d.o.f.}$ $91/57$ $109/91$
$f_{\mathrm{2-10}}\,\mathrm{(erg\,cm^{-2}\,s^{-1})}$ $(1.2\pm 0.1)\times 10^{-10}$ $6.4^{+0.3}_{-0.2}\times 10^{-11}$
footnotetext: Radius calculated assuming a distance of $8$ kpc, without correcting for scattering in the neutron star atmosphere.
We analyzed the spectra using Xspec (v12.14.1 Arnaud, 1996). In order to perform parameter estimation, we used the $W$ fit statistic (Wachter et al., 1979), which is used by Xspec in the case of Poisson source and background spectra. We used the Chi-square test statistic to assess goodness-of-fit. For both observations, the number of background counts becomes comparable to the number of source counts above around $30$ keV. Therefore we only modeled the spectra for energies in the range 3–30 keV. Throughout our analysis, we performed simultaneous fits to the FPMA and FPMB spectra, applying a multiplicative constant to account for small differences in the overall normalization between the two spectra.
Due to the lack of evidence for strong absorption and to the insensitivity to absorption in the NuSTAR band, for all spectral models we applied an absorption column with fixed density using the multiplicative Xspec model component, tbabs. We fixed the column density to a value of $N_{\mathrm{H}}=10^{21}\,\mathrm{cm^{-2}}$, determined using the full-sky HI survey, HI4PI (HI4PI Collaboration et al., 2016). We first attempted to model the NuSTAR spectra using a simple absorbed power law with a high-energy cutoff. We found that this did not produce an adequate fit, resulting in a test statistic of $\chi^{2}/\mathrm{d.o.f.}=142/58$ for the first observation and $\chi^{2}/\mathrm{d.o.f.}=286/92$ for the second observation.
Given the superburst interpretation presented in Aoyama et al. (2025), we added a physically-motivated single-temperature blackbody component to our spectral models, representing thermal emission from the surface of a neutron star heated by either thermonuclear burning during the burst or by persistent accretion flow onto a boundary layer. Notably, the peak flux observed by MAXI/GSC during the burst was around $2\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. Although the distance to MAXI J1752--457 is unknown, the source is located at $(\alpha,\delta)=(268.269,-45.866)$, or $(l,b)=(346.092,-9.8517)$, placing it at a $17^{\circ}$ separation from the Galactic center. Like Aoyama et al. (2025), we tentatively assume that the angular proximity of the source to the Galactic center translates into physical proximity, giving a distance of around 8 kpc. This assumption leads to a derived peak luminosity on the order of $10^{38}$$\mathrm{erg\,s^{-1}}$, very near the Eddington luminosity for spherical accretion onto a neutron star: $L_{\mathrm{Edd}}=1.38\left(M/{M_{\odot}}\right)\times 10^{38}$$\mathrm{erg\,s^{-1}}$, where the canonical neutron star mass is $M=1.4M_{\odot}$. Thus, the physical interpretation of the source as a bursting neutron star is consistent with a distance of around 8 kpc, and we therefore present derived luminosities and blackbody radii using this fiducial value.
We found that when we included a single-temperature blackbody component, the value of the high-energy cutoff was pushed far beyond the energy range we considered, indicating that the cutoff is not necessary to fit the data. The combination of a single-temperature blackbody and a simple power law component fits the data significantly better than a cutoff power law alone, resulting in a test statistic of $\chi^{2}/\mathrm{d.o.f.}=91/57$ for the first observation and $\chi^{2}/\mathrm{d.o.f.}=109/91$ for the second observation. We found that the spectra can be described without significant changes in the blackbody temperature and the spectral index between the two observations. Instead, the changes between the spectra can be almost entirely described by changes in overall flux, with both the blackbody flux (i.e. radius) and power law flux decreasing between the first and second observations. We measured a decrease in unabsorbed flux of almost one half from $(1.2\pm 0.1)\times 10^{-10}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$during the first observation to $(6.4^{+0.3}_{-0.2})\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$during the second observation, corresponding to luminosities of around $9\times 10^{35}$$\mathrm{erg\,s^{-1}}$and $5\times 10^{35}$$\mathrm{erg\,s^{-1}}$.
In Table 2, we list the resulting blackbody and power law parameters, the test statistic achieved with the model, and the unabsorbed 2–10 keV flux measured for both observations. In Figure 3 we show the FPMA and FPMB spectra, with the first observation shown on the left and the second observation shown on the right. We also plot the best-fit model including the single-temperature blackbody and the power law components. In addition to the good test statistic achieved with this model for the second observation, we do not observe residual features which would indicate that an additional component is necessary to describe the spectra.
We also considered slightly different forms of this basic spectral model. For example, we replaced the single-temperature blackbody with a multi-temperature disk blackbody. This model resulted in a comparable but slightly worse test statistic. This model yields an inner disk temperature of around $kT_{\mathrm{in}}=0.7$ keV and a slightly shallower power law with a photon index of $\Gamma=3.4$. Similarly, replacing the simple power law component with a thermally Comptonized continuum component (nthcomp; Zdziarski et al., 1996; Życki et al., 1999), did not improve the test statistic. In this case, the relevant spectral parameters did not change compared to the simple power law model.
In order to check the self-consistency of the spectral model, we also reproduced the spectral fits using simpl (Steiner et al., 2009) rather than an analytic power law. We again found no appreciable change in the test statistic. This model yields a slightly larger blackbody radius (at 8 kpc, uncorrected for atmospheric scattering effects) of $11^{+2}_{-1}$ km and $7.4^{+0.6}_{-0.5}$ km for the first and second NuSTAR observations, respectively. For both observations, using simpl instead of an analytic power law results in a decrease in the blackbody temperature by a few percent, an increase in the photon index of about 0.1, and an upscattering fraction of around $10\%$, assuming that the power law is seeded by the spherical blackbody component.
We note that the test statistic is comparably poor for the first observation, leading to a null-hypothesis probability of $0.3\%$. We find that the residuals for these spectra are not well matched between FPMA and FPMB (see the bottom left panel of Figure 3). We therefore hypothesize that the higher test statistic is primarily due to systematic error introduced by the complicated Mode 06 extraction procedure with source regions bordering the stray light region. While the inferred parameters provide a point of comparison with the second observation, we do not rely heavily on the first observation to inform our conclusions.
3.2 Timing analysis
Figure 4: Left: Leahy-normalized power density spectrum (PDS) of the 3–30 keV NuSTAR source (black) and background (magenta) light-curves measured during the second observation. The PDS have been binned logarithmically for clarity, and the unbinned source PDS is shown in gray. The only clear feature we observe is a low-frequency noise component, which we fit with a power law, shown in red. Right: Leahy-normalized source PDS calculated for the 3–6 keV (green) and 6–30 keV (orange) light-curves. The low-frequency component is somewhat more pronounced in the hard X-ray band. The effects of deadtime were corrected using the Fourier Amplitude Difference (FAD) normalization technique.
Table 3: Details of the NuSTAR timing analysis.
Energy band Power law index Source rms (%)
(keV) ($10^{-3}$$10^{-1}$ Hz)
3–30 $1.8^{+0.3}_{-0.2}$ $17\pm 3$
3–6 $1.8^{+0.7}_{-0.4}$ $13\pm 4$
6–30 $2.0^{+0.4}_{-0.3}$ $58\pm 8$
We investigated the timing properties of MAXI J1752--457 during the second NuSTAR observation. Beginning with the clean Mode 1 and Mode 06 event files, we first corrected the photon arrival times for the motion of NuSTAR about the barycenter of the Solar System using the ftools utility, barycorr. We specified the source position as measured by Swift/XRT and reported by Liu et al. (2024). We used NuSTAR clockfile 20100101v203, produced on April 15, 2025. Next, we filtered the events using circular source regions with radii $100^{\prime\prime}$ in order to encompass as many source counts as possible given the centroid smearing in Mode 06. Unlike our spectral analysis, and in order to avoid complicated good time intervals (GTIs), we did not split the Mode 06 events into different CHU combinations.
We next used the timing analysis software packages HENDRICS and Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024a, b) to complete the rest of the timing analysis. We found that the background dominates the source spectrum above 30 keV, so we filtered out events with energy below 3 keV and above 30 keV. For each focal plane module, we merged the Mode 1 and Mode 06 event lists. Using Stingray, we converted the event lists into light-curves with bins of $dt=2048^{-1}$ s. We split these light-curves into 16 segments of length $1024$ s each. We calculated the averaged power density spectrum (PDS) over these segments for FPMA and FPMB, then produced a combined PDS which we corrected for deadtime using the Fourier Amplitude Difference normalization described in Bachetti & Huppenkothen (2018). We repeated this procedure in order to produce a background PDS, using only events in background regions with radius $150^{\prime\prime}$, away from the source. The resulting Leahy-normalized PDS, binned logarithmically for clarity, are shown in the left panel of Figure 4, where the source PDS is shown in black, and the background PDS is shown in magenta. In order to investigate the energy-dependence of the source variability, we also calculated the source PDS for events with 3–6 keV and for events with 6–30 keV, shown in green and orange, respectively, in the right panel of Figure 4.
The PDS is relatively featureless aside from a clear low-frequency noise component. We do not find any evidence for pulsations or quasiperiodic oscillations (QPOs). We performed a least-squares fit of the PDS with a power law model. We also fit this model to the “soft” (3–6 keV) and “hard” (6–30 keV) PDS. The best-fit power law index for each energy band is listed in Table 3, along with the fractional rms calculated for the frequency range $10^{-3}$$10^{-1}$ Hz. Our results indicate that the low-frequency noise may be slightly steeper in the hard band, but the power law index is consistent with 2 within the $90\%$ confidence intervals regardless of the energy band: for the 3–6 keV range, we find a power law index of $1.8^{+0.3}_{-0.2}$, and for the 6–30 keV range, we find a power law index of $2.0^{+0.4}_{-0.3}$. Notably, the low frequency noise is significantly higher in the hard band than in the soft band.
4 Discussion and Conclusions
MAXI and NinjaSat observations of MAXI J1752--457 during a superburst in November, 2024, showed that the flux and spectral evolution were consistent with an accretion-induced superburst (Aoyama et al., 2025). While these observations were sensitive to the blackbody component of the spectrum, and have demonstrated clear evolution of the temperature and radius of the blackbody, NuSTAR’s unique sensitivity above $10$ keV has allowed us to put far better constraints on the blackbody parameters following the rapid decay of the burst. We found that by the time of the second NuSTAR observation, which occured about $79$ hours after the onset of the burst, the blackbody temperature had decreased to $0.60\pm 0.1$ keV, and the inferred radius (assuming a distance of $8$ kpc) had shrunk to $6.0^{+0.4}_{-0.3}$ km.
Figure 5: The change in $W$-statistic with respect to the best-fit value when modeling the spectra from the second NuSTAR observation with an absorbed blackbody and cutoff power law. $\Delta W$ is plotted as a function of the power law photon index, $\Gamma$, and cutoff energy, $E_{\mathrm{cut}}$. The $68\%$ (1-$\sigma$), $90\%$, and $99\%$ confidence regions for these parameters are shown as black, purple, and white contours, respectively. The dashed red line marks the best-fit value, $\Gamma=3.7$, when there is no high-energy cutoff. When the cutoff is included, $E_{\mathrm{cut}}$ is unconstrained, with a $99\%$ lower limit of 8 keV.
NuSTAR observations also revealed that, in addition to a blackbody component, the source spectrum contained a hard X-ray excess which is well-described by a steep power law with photon index $\Gamma\approx 4$. This indicates the presence of a non-thermal plasma in which soft X-ray photons gain energy via Compton upscattering. Accreting neutron stars at similar luminosities ($10^{35-36}$$\mathrm{erg\,s^{-1}}$) tend to exhibit power law spectral components with photon indices in the range $3.01.5<\Gamma<$ 3.0 (Wijnands et al., 2015). We therefore revisit the spectra obtained during the second NuSTAR observation in order to explore the possibility of a more typical photon index.
As we noted in Section 3.1, when we allowed for a high-energy cutoff to be applied to the power law component, the best-fit value of $E_{\mathrm{cut}}$ was far beyond the NuSTAR energy range, indicating that a power law without a cutoff energy is preferred. However, including a high-energy cutoff allows for more flexibility in the photon index parameter as a low cutoff energy can compensate for a shallower power law. For the cutoff power law model, the $90\%$ confidence interval for the photon index is expanded, giving $\Gamma=3.7^{+0.2}_{-0.7}$. Furthermore, we calculated the $99\%$ confidence intervals for the photon index and cutoff energy by varying each parameter independently until the fit statistic had increased by 6.635. We found intervals of $4.02.3<\Gamma<$ 4.0 and $E_{\mathrm{cut}}>8$ keV. In order to explore the parameter space in even more detail, we stepped over the photon index and the cutoff energy and calculated the $W$-statistic at each step. The resulting grid is shown in Figure 5, with contours representing the 1-$\sigma$, $90\%$, and $99\%$ confidence regions. We again find that the spectra strongly prefer a steep power law with a $90\%$ confidence region that indicates a photon index greater than 2.5 and a cutoff energy above 10 keV. While photon index values of $<2.5$ are allowed within the 99% confidence region, they would require cutoff energies of less than 10 keV. This combination of parameters would imply a very high optical depth (Kara & García, 2025). While we cannot completely rule out a more typical value for the photon index, spectral modeling of MAXI J1752--457 strongly prefers an abnormally steep power law in order account for the hard X-ray excess above about 7 keV.
Existing analyses of the hard X-ray spectra of accreting neutron stars during and after superburst are limited due to the relative infrequency of these bursts. In fact, at the time of writing, we are unable to identify any published observations of superbursts with NuSTAR. However, the Rossi X-ray Timing Explorer (RXTE), thanks to its All-Sky Monitor (ASM) instrument, observed 12 superbursts in the hard X-ray during its lifetime (in’t Zand, 2017). Often, analyses of the spectral evolution of superbursts with RXTE fixed the power law component to the value measured for persistent emission, typically around $\Gamma\approx 2$ (e.g. Keek et al., 2008). In some instances, significant softening of the superburst decay spectrum compared to the persistent spectrum has been observed (e.g. Kuulkers et al., 2004).
Studies of superbursts in the soft X-ray band ($<10$ keV), although insensitive to the hard X-ray emission, have hinted at high photon indices during or shortly after superbursts or similar bursts. For example, Bult et al. (2021) reported on NICER measurements of the spectral evolution of IGR J17062--6143 during an intermediate duration Type I X-ray burst observed in 2020. While not categorized as a superburst, the decay of this burst lasted several days, making it comparable to the case of MAXI J1752--457. These authors reported that, when a relativistic reflection component was included in the spectral model of the decay emission, the photon index reached values in the range of $53<\Gamma<$ 5 throughout the two-day period following the burst before settling to more typical values around $\Gamma\approx 2.5$. Notably, however, when the spectrum was modeled without relativistic reflection, the opposite trend in the power law photon index was recovered, with $\Gamma$ increasing from around 1 to 2 throughout the burst decay.
In a recently published NICER study of the 2021 superburst of 4U 1820--30, Peng et al. (2025) modeled the persistent, burst, and decay emission using a combination of a spherical blackbody and a Comptonization component represented by the Xspec model compTT. The authors noted that during the burst, the contribution of the Comptonization component could be neglected, but they reported that during the first few hours of the decay, they measured a significantly lower electron temperature than during the persistent emission, indicating a softened hard X-ray component. Both of these NICER studies, however, highlight the difficulty of disentangling spectral components with soft X-ray observations alone when the emission is dominated by a bright blackbody component.
In light of the hard and soft X-ray analyses of superburst and decay spectral evolution which have been reported thus far, we conclude that NuSTAR observations of a photon index of $\Gamma\approx 4$ following the 2024 superburst of MAXI J1752--457 represents a new insight into the interaction of superburst emission with the corona. While it has been shown that Type I X-ray bursts result in cooling corona (Speicher et al., 2020), superbursts offer an opportunity to study more extended interaction between burst emission and the corona. Although we did not measure the persistent emission before the burst in order to compare to our observations, given that the photon index we measured is so atypical of persistent LMXB emission, our results indicate that changes to the corona may last on the order of days after a superburst.
Our analysis of the source variability did not reveal discrete features like pulsations or quasiperiodic oscillations (QPOs). Instead, we found that the source variability was dominated by red noise, which we modeled with a power law. We found a power law index of $1.8^{+0.3}_{-0.2}$, which is typical for the low-frequency noise seen in XRBs (Belloni et al., 2002), and we found that the low-frequency variability is more pronounced at higher energies. Using the Xspec tool, cpflux, we determined that the spectral power law component makes up $22\%$ of the source photon flux in the 3–6 keV range and $77\%$ of the source photon flux in the 6–30 keV range. Given a total source rms of $(13\pm 4)\%$ and $(58\pm 8)\%$ in the soft and hard bands, respectively, and if we assume that the Comptonization component is the sole source of red noise (i.e., the rms variability of the blackbody component is very small), we infer that the $10^{-3}-10^{-1}$ Hz rms of the Comptonization component was $(59\pm 18)\%$ in the 3–6 keV range and $(75\pm 10)\%$ in the 6–30 keV range. In other words, we find that the energy-dependent photon flux fraction of the spectral power law component is consistent with our measurements of the source rms as a function of energy under the assumption that the non-thermal component is the primary source of the low-frequency variability.
The significant difference in the degree of variability between the thermal and power law components raises the question of the origin of the X-ray emission during the NuSTAR observations. The thermal emission may originate from radiative cooling of the neutron star surface following the superburst, or it may be primarily due to a boundary layer heated by disk accretion (Popham & Sunyaev, 2001). Low-frequency variability is typically attributed to variable thermal disk emission for both black hole and neutron star sources (Titarchuk et al., 2007). Given that the low-frequency noise we observe can be attributed primarily to the power law spectral component, rather than to the thermal component, this would suggest that the corona is seeded by soft X-ray photons from an accretion disk such that disk variability propagates through the Comptonizing medium. Several other pieces of evidence point towards accretion powered emission. We do not observe clear evidence for cooling of the blackbody component between the two NuSTAR observations, and we find that the flux ratio between the blackbody component and the power law component did not change significantly between the two NuSTAR observations, suggesting that both the thermal and power law components were powered by accretion such that the change in accretion rate affected the overall normalization of the spectra but not the spectral shape. Furthermore, by the time that the NuSTAR observations were carried out, the source flux had become comparable to that observed by MAXI/GSC prior to the superburst. In other words, while we cannot rule out the radiative cooling scenario, we favor the interpretation that the hard X-ray emission observed by NuSTAR several days after the superburst peak was due primarily to disk accretion.
These measurements demonstrate the importance of rapid follow-up of transient sources with focusing hard X-ray observatories like NuSTAR in order to place superior constraints on the spectral evolution and physical properties of bursting sources. Further hard X-ray observations of superbursts will help to elucidate the behavior of the corona before, during, and after these long-lasting thermonuclear flashes.
This work was partially supported under NASA Grant No. 80NSSC25K7455 and made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. We thank the NuSTAR Operations, Software, and Calibration teams for support with the execution and analysis of these observations. This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS), jointly developed by the ASI Science Data Center (ASDC, Italy) and the California Institute of Technology (USA). This research has also made use of the MAXI data provided by RIKEN, JAXA and the MAXI team.